The value of a share of a company can be represented by V(x) = x2 – 6x + 13, where x is the number of months after January 2004. What is the lowest value V(x) will reach, and when did that occur?
Answer by stanbon(75887) 
 The value of a share of a company can be represented by V(x) = x2 – 6x + 13, where x is the number of months after January 2004. What is the lowest value V(x) will reach, and when did that occur?
---
V is a quadratic with a=1,b=-6,c = 13
---
Mimum V occurs at x = -b/2a = 6/2 = 3
----------------------------------
The minimum V value = V(3) = 3^2-6*3+13 = 16
=============================================
Cheers,
Stan H.
|
|
|